Which of the following transitions (in a hydrogen atom) represents an emission of the longest wavelength photon?

a) n = 1 to n = 2

b) n = 4 to n = 2

c) n = 3 to n = 1

d) n = 5 to n = 4

e) n = 3 to n = 5


1
Expert's answer
2021-10-05T01:33:40-0400

Solution:

Rydberg equation can be used:

1/λ = RH × (1/n2f - 1/n2i)

where λ is the wavelength of the emitted photon, RH is the Rydberg constant (RH=1.097×107 m−1), and nf and ni are the final and initial energy states, respectively.


For an emission to occur, the hydrogen atom should move from a higher energy state to a lower energy state. In this case, choices (a) and (e) are incorrect since they indicate the excitation of the hydrogen atom from a lower to a higher energy level.


Thus:

(b) ni = 4 to nf = 2

1/λ = (1.097×107 m−1) × (1/22 - 1/42) = 2056875

λ = 1 / 2056875 = 4.86×10−7 m−1 = 486 nm


(c) ni = 3 to nf = 1

1/λ = (1.097×107 m−1) × (1/12 - 1/32) = 9751111.111

λ = 1 / 9751111.111 = 1.0255×10−7 m−1 = 102.55 nm


(d) ni = 5 to nf = 4

1/λ = (1.097×107 m−1) × (1/42 - 1/52) = 246825

λ = 1 / 246825 = 405.1×10−6 m−1 = 4051 nm


102.55 nm (c) < 486 nm (b) < 4051 nm (d)


Based on the obtained values, the transition that represents the emission of the longest wavelength photon is n = 5 to n = 4 (i.e., 4050 nm).


Answer: d) n = 5 to n = 4